How many protons, neutrons, and electrons are present in one Ti+2 ion with an atomic weight (mass number) of 46?
1
Expert's answer
2017-02-27T15:39:05-0500
Atomic number = 22
p = 22
e ̅=22-2=20
n = [Mass number]-[Atomic number]=46-22=24